OpenStudy (anonymous):

solve by completing the square: x^2-10x-18=0

OpenStudy (anonymous):

x^2 - 10x + 25 - 25 - 18 = 0 x^2 - 10x + 25 = 43 (x-5)^2 = 43 x-5 = SQRT(43) x = SQRT(43) + 5 You get two values because SQRT(43) can be positive or negative

OpenStudy (anonymous):

-6.56 + 5 +6.56 + 5

OpenStudy (anonymous):

GT where you from man?

OpenStudy (anonymous):

I simply don't like the representation used by kumar2006 because SQRT(43) is written as is and that in itself will result in a +ve and -ve values.

OpenStudy (anonymous):

Funny.....I would have written it as 5 + SQRT(43). When you take the square root of 43, you get positive and negative 6.56.